# Achieving double-logarithmic precision dependence in optimization-based quantum unstructured search

### 27 Mar 2026 v1

@@ Abstract
Grover’s algorithm is a fundamental quantum algorithm that achieves a quadratic speedup for unstructured search problems of size $N$. Recent studies have reformulated this task as a maximization problem on the unitary manifold and solved it via linearly convergent Riemannian gradient ascent (RGA) methods, resulting in a complexity of $\mathcal{O}(\sqrt{N}\log(1/\varepsilon))$. In this work, we adopt the Riemannian modified Newton (RMN) method to solve the quantum search problem. We show that, in the setting of quantum search, the Riemannian Newton direction is collinear with the Riemannian gradient in the sense that the Riemannian gradient is always an eigenvector of the corresponding Riemannian Hessian. As a result, without additional overhead, the proposed RMN method numerically achieves a quadratic convergence rate with respect to error $\varepsilon$, implying a complexity of $\mathcal{O}(\sqrt{N}\log\log(1/\varepsilon))$, which is double-logarithmic in precision. Furthermore, our approach remains Grover-compatible, namely, it relies exclusively on the standard Grover oracle and diffusion operators to ensure algorithmic implementability, and its parameter update process can be efficiently precomputed on classical computers.
@@ I Introduction
Unstructured search, the task of finding specific target elements in an unsorted space of size $N$, is a ubiquitous challenge across cryptographic analysis [17, 5], optimization [13, 15], and machine learning [11, 12]. Classically, this problem is inherently inefficient, as guaranteeing success requires $\Omega(N)$ queries [4]. This constitutes a severe computational bottleneck for large-scale data retrieval [28, 40].
Grover’s quantum search algorithm resolved this bottleneck by achieving an optimal query complexity of $\Theta(\sqrt{N})$ [18, 43, 54, 3], marking a landmark demonstration of quantum advantage. Beyond its direct application to the search problem, Grover’s mechanism is a fundamental primitive in quantum computing. Its principles form the basis of amplitude amplification [7, 47] and have been further generalized within the broader framework of quantum singular value transformation (QSVT) [16, 41].
@@
Specifically, a recent study [30] reformulates Grover’s search as a maximization problem on the unitary manifold and proposes a physically implementable Riemannian gradient ascent (RGA) method to achieve the desired quadratic speedup in size $N$. While this first-order approach establishes a rigorous link between the optimization framework and quantum dynamics, it is limited to linear convergence with respect to the accuracy parameter $\varepsilon$, with total complexity $\mathcal{O}(\sqrt{N}\log(1/\varepsilon))$.
To further accelerate convergence in terms of the target accuracy $\varepsilon$, second-order optimization methods, such as the Riemannian Newton method [2, 1], provide a natural approach. In this work, we propose a tailored Riemannian modified Newton (RMN) method for Grover’s search problem, where each iteration admits a physically implementable second-order update using only the oracle and diffusion operators, analogous to Grover’s algorithm.
@@
We test our RMN method numerically and find that, once the iterates get close to the target state, RMN can attain very high-precision solutions in very few iterations. Furthermore, our numerical results show that RMN achieves a quadratic convergence rate with respect to the accuracy $\varepsilon$, with total complexity $\mathcal{O}(\sqrt{N}\log\log(1/\varepsilon))$. Our work advances the optimization-based framework for quantum algorithm design.
@@ II.2 Manifold optimization formulation
When discussing the tangent space $T_{U}=\mathfrak{u}(N)U$, we often focus on $\mathfrak{u}(N)$ itself and omit the right unitary $U$. [30, Theorem 3.2] has shown that $\widetilde{\operatorname{grad}}f\left(U^{\star}\right)=0$ if and only if $U^{\star}\in\mathrm{U}(N)$ globally minimizes $f(U)$ over $\mathrm{U}(N)$ with the minimum value 0, or globally maximizes it with the maximum value 1.
@@ III.2 Grover-compatible retractions
Theorem 1 (Invariant 2D gradient subspace [30, Theorem 3.3]).
@@
In fact, retractions satisfying the above definition are easy to construct, and their forms may vary. Here we introduce a $5$-factor retraction in Example˜1. According to the number $K$ of factors of the form $e^{i\theta H}$ and $e^{i\theta\psi_{0}}$ it contains, we refer to it as a $K$-factor retraction. More examples, such as $6$-factor and $8$-factor retractions, can be found in [30, Proposition 3.5].
@@ 18  Set $k:=k+1$;
[30, Theorem 4.6] provides explicit results on the convergence and complexity of Algorithm˜1. Suppose we run Algorithm˜1 with the 5-factor retraction in Eq.˜24, and choose a fixed step size $t_{k}=1/L_{\text{Rie}}$, where $L_{\text{Rie}}:=2+N/\sqrt{2M(N-M)}\in\mathcal{O}(\sqrt{N/M})$. Then, for any $0<\varepsilon\leq q_{0}$, the iterates satisfy $1-q_{T}\leq\varepsilon$ within at most $T=\left\lceil 6L_{\text{Rie}}\log\left(1/\varepsilon\right)\right\rceil$ iterations. In summary, we establish an iteration complexity of $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$ to reach an $\varepsilon$-global maximizer, which is consistent with Grover’s quadratic speedup. The term $\log(1/\varepsilon)$ indicates that the method converges linearly with respect to the accuracy $\varepsilon$.
@@ III.3 Classical simulability of Algorithm˜1
Algorithm˜1 does not need to be executed directly on a quantum device. As shown in Theorem˜2, the procedure is classically simulable. Intuitively, the entire evolution of Algorithm˜1 takes place within a two-dimensional subspace. We choose $u:=H\left|\psi_{0}\right\rangle$ and $v:=(I-H)\left|\psi_{0}\right\rangle$ as a basis for the Grover plane. It therefore suffices to represent the state $\left|\psi_{k}\right\rangle$ and the gradient $\left[H,\psi_{k}\right]$ in this basis. The detailed proof can be found in [30, Theorem 3.8]. In practice, the gate angles can be computed in advance on a classical computer, after which the resulting circuit is executed on quantum hardware.
Theorem 2 (Classical simulability of Algorithm˜1 [30, Theorem 3.8]).
@@ IV.1 Riemannian Hessian
The self-adjointness means that $\langle\Xi,\widetilde{\operatorname{Hess}}f(U)[\Omega]\rangle=\langle\widetilde{\operatorname{Hess}}f(U)[\Xi],\Omega\rangle$ for any $\Omega,\Xi\in\mathfrak{u}(N)$. A derivation of this Hessian expression can be found in [31, Proposition 1]. An equivalent definition is based on the second-order Taylor expansion on the manifold:
@@ IV.2 Grover-compatible Riemannian modified Newton method
After ensuring that the Newton direction $\Omega_{k}^{\mathrm{N}}$ is an ascent direction, we further employ an Armijo backtracking line search; see [44, Algorithm 3.2]. Specifically, we seek the smallest integer $m\geq 0$ such that $t_{k}=\rho^{m}$ satisfies
@@
where $\rho=0.5$ and $c=10^{-4}$. Note that $\left\langle\Omega_{k}^{\mathrm{N}},\mathrm{g}_{k}\right\rangle>0$, which guarantees that the above procedure terminates after finitely many iterations; see [44, Lemma 3.1]. Consequently, the cost value increases monotonically. Finally, a complete description of the proposed algorithm is given in Algorithm˜2.
This modified Newton method with line search ensures convergence from any initial point (i.e., it enjoys the global convergence property of first-order gradient methods), while achieving quadratic convergence once the iterate is sufficiently close to the solution. Intuitively, during the early iterations, Algorithm˜2 behaves similarly to Algorithm˜1. When the iterate is sufficiently close to the solution (so that $q_{k}>0.5$ and no direction modification is required), the Armijo condition is always satisfied with $t_{k}=1$; see [44, Theorem 3.6 & 3.8]. Consequently, Algorithm˜2 reduces to the standard Newton method.
@@ V Numerical experiments
RGA is shown in [30] to have complexity $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$. Fixing the accuracy $\varepsilon$ and setting $M=1$, this reduces to the well-known quadratic speedup $\mathcal{O}(\sqrt{N})$, i.e., the iteration count grows linearly with $\sqrt{N}$. The proposed RMN is a second-order optimization method, whose improvement targets the dependence on the accuracy $\varepsilon$ rather than on the problem size NN [^3] Optimization literature typically fixes the problem size and studies how to achieve higher accuracy with fewer iterations.. It is therefore expected to retain the same scaling $\mathcal{O}(\sqrt{N})$ with respect to the problem size. However, due to its quadratic convergence, the overall complexity improves to $\mathcal{O}(\sqrt{N/M}\log\log(1/\varepsilon)).$
@@ VI Discussion
In this work, we develop a deeper understanding of the quantum search problem through the lens of the manifold optimization framework by introducing the Grover-compatible Riemannian modified Newton (RMN) method. While the previous first-order Riemannian gradient ascent (RGA) method [30] successfully recasts quantum search as a maximization problem on the unitary group and achieves $\mathcal{O}(\sqrt{N}\log(1/\varepsilon))$ query complexity, it is limited to a linear convergence rate with respect to the target accuracy $\varepsilon$. By incorporating second-order geometric information, we overcome this limitation.
The cornerstone of our RMN method is the theoretical observation that, for projector Hamiltonians $H=H^{2}$, the Riemannian gradient is an eigenvector of the corresponding Riemannian Hessian. This property eliminates the need for costly Hessian inversion, reducing the Newton update to a scaled gradient step. Consequently, our RMN method can achieve a quadratic convergence rate, driving the total complexity down to $\mathcal{O}(\sqrt{N}\log\log(1/\varepsilon))$ without incurring additional overhead per iteration. Furthermore, we demonstrate that the classical simulability of the parameter update process is also preserved, allowing for the efficient precomputation of gate angles prior to execution on quantum hardware.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China under the grant numbers 12501419, 12331010 and 12288101, the National Key R&D Program of China under the grant number 2024YFA1012901, the Quantum Science and Technology-National Science and Technology Major Project via Project 2024ZD0301900, and the Fundamental Research Funds for the Central Universities, Peking University.
@@ A.4 Riemannian Hessian
A derivation of this Hessian expression can be found in [31, Proposition 1]. To verify the correctness of Eq.˜60, we will apply the second-order Taylor expansion on $\mathrm{U}(N)$. For $\eta=\Omega U\in T_{U}$, using the Riemannian exponential retraction $\mathrm{R}_{U}(\eta)=e^{\Omega}U$ in Eq.˜54, we obtain

### 03 Apr 2026 v2

@@ Abstract
Grover’s algorithm is a fundamental quantum algorithm that achieves a quadratic speedup for unstructured search problems of size $N$. Recent studies have reformulated this task as a maximization problem on the unitary manifold and solved it via linearly convergent Riemannian gradient ascent (RGA) methods, resulting in a complexity of $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$, where $M$ denotes the number of target items. In this work, we adopt the Riemannian modified Newton (RMN) method to solve the quantum search problem, under the assumption that the ratio $M/N$ is known. We show that, in this setting, the Riemannian Newton direction is collinear with the Riemannian gradient in the sense that the Riemannian gradient is always an eigenvector of the corresponding Riemannian Hessian. As a result, without additional overhead, the proposed RMN method numerically achieves a quadratic convergence rate with respect to the error $\varepsilon$, implying a complexity of $\mathcal{O}(\sqrt{N/M}\log\log(1/\varepsilon))$. Furthermore, our approach remains Grover-compatible, namely, it relies exclusively on the standard Grover diffusion and oracle operators to ensure algorithmic implementability, and its parameter update process can be efficiently precomputed on classical computers.
@@ I Introduction
Unstructured search, the task of finding specific target elements in an unsorted space of size $N$, is a ubiquitous challenge across cryptographic analysis [17, 5], optimization [13, 15], and machine learning [11, 12]. Let $M\ll N$ denote the number of target items. Classically, this problem is inherently inefficient, as guaranteeing success requires $\Omega(N/M)$ queries [4]. This constitutes a severe computational bottleneck for large-scale data retrieval [29, 42].
Grover’s quantum search algorithm resolved this bottleneck by achieving an optimal query complexity of $\Theta(\sqrt{N/M})$ [18, 45, 56, 3], marking a landmark demonstration of quantum advantage. Beyond its direct application to the search problem, Grover’s mechanism is a fundamental primitive in quantum computing. Its principles form the basis of amplitude amplification [7, 49] and have been further generalized within the broader framework of quantum singular value transformation (QSVT) [16, 43].
@@
Specifically, a recent study [31] reformulates Grover’s search as a maximization problem on the unitary manifold and proposes a physically implementable Riemannian gradient ascent (RGA) method to achieve the desired quadratic speedup in size $N$. While this first-order approach establishes a rigorous link between the optimization framework and quantum dynamics, it is limited to linear convergence with respect to the accuracy parameter $\varepsilon$, with total complexity $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$.
To further accelerate convergence in terms of the target accuracy $\varepsilon$, second-order optimization methods, such as the Riemannian Newton method [2, 1], provide a natural approach. In this work, we propose a Riemannian modified Newton (RMN) method for Grover’s search problem, where each iteration admits a physically implementable second-order update using only the diffusion and oracle operators, analogous to Grover’s algorithm.
@@
We test our RMN method numerically and find that, once the iterates get close to the target state, RMN can attain very high-precision solutions in very few iterations. Furthermore, our numerical results show that RMN achieves a quadratic convergence rate with respect to the accuracy $\varepsilon$, with total complexity $\mathcal{O}(\sqrt{N/M}\log\log(1/\varepsilon))$. However, we emphasize that, in the RMN method, the rotation angle at each iteration must be precomputed on a classical computer under the assumption that the exact value of the ratio $\tau=M/N$ is known. Indeed, when $\tau$ is known, there exist variants of Grover’s algorithm [21, 37, 7] that achieve success probability exactly equal to $1$, with complexity independent of $\varepsilon$. Hence, the primary contribution of this work is not to improve the practical complexity over the existing search algorithms, but rather to provide a new optimization-based method and perspective for quantum unstructured search.
@@ II.2 Manifold optimization formulation
When discussing the tangent space $T_{U}=\mathfrak{u}(N)U$, we often focus on $\mathfrak{u}(N)$ itself and omit the right unitary $U$. [31, Theorem 3.2] has shown that $\widetilde{\operatorname{grad}}f\left(U^{\star}\right)=0$ if and only if $U^{\star}\in\mathrm{U}(N)$ globally minimizes $f(U)$ over $\mathrm{U}(N)$ with the minimum value 0, or globally maximizes it with the maximum value 1.
@@ III.2 Grover-compatible retractions
Theorem 1 (Invariant 2D gradient subspace [31, Theorem 3.3]).
@@
In fact, retractions satisfying the above definition are easy to construct, and their forms may vary. Here we introduce a $5$-factor retraction in Example˜1. According to the number $K$ of factors of the form $e^{i\theta H}$ and $e^{i\theta\psi_{0}}$ it contains, we refer to it as a $K$-factor retraction. More examples, such as $6$-factor and $8$-factor retractions, can be found in [31, Proposition 3.5].
@@ 18  Set $k:=k+1$;
[31, Theorem 4.6] provides explicit results on the convergence and complexity of Algorithm˜1. Suppose we run Algorithm˜1 with the 5-factor retraction in Eq.˜24, and choose a fixed step size $t_{k}=1/L_{\text{Rie}}$, where $L_{\text{Rie}}:=2+N/\sqrt{2M(N-M)}\in\mathcal{O}(\sqrt{N/M})$. Then, for any $0<\varepsilon\leq q_{0}$, the iterates satisfy $1-q_{T}\leq\varepsilon$ within at most $T=\left\lceil 6L_{\text{Rie}}\log\left(1/\varepsilon\right)\right\rceil$ iterations. In summary, we establish an iteration complexity of $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$ to reach an $\varepsilon$-global maximizer, which is consistent with Grover’s quadratic speedup. The term $\log(1/\varepsilon)$ indicates that the method converges linearly with respect to the accuracy $\varepsilon$.
@@ III.3 Classical simulability of Algorithm˜1
Algorithm˜1 does not need to be executed directly on a quantum device. As shown in Theorem˜2, the procedure is classically simulable. Intuitively, the entire evolution of Algorithm˜1 takes place within a two-dimensional subspace. We choose $u:=H\left|\psi_{0}\right\rangle$ and $v:=(I-H)\left|\psi_{0}\right\rangle$ as a basis for the Grover plane. It therefore suffices to represent the state $\left|\psi_{k}\right\rangle$ and the gradient $\left[H,\psi_{k}\right]$ in this basis. The detailed proof can be found in [31, Theorem 3.8]. In practice, the gate angles can be computed in advance on a classical computer, after which the resulting circuit is executed on quantum hardware.
Theorem 2 (Classical simulability of Algorithm˜1 [31, Theorem 3.8]).
@@ IV.1 Riemannian Hessian
The self-adjointness means that $\langle\Xi,\widetilde{\operatorname{Hess}}f(U)[\Omega]\rangle=\langle\widetilde{\operatorname{Hess}}f(U)[\Xi],\Omega\rangle$ for any $\Omega,\Xi\in\mathfrak{u}(N)$. A derivation of this Hessian expression can be found in [32, Proposition 1]. An equivalent definition is based on the second-order Taylor expansion on the manifold:
@@ IV.2 Grover-compatible Riemannian modified Newton method
After ensuring that the Newton direction $\Omega_{k}^{\mathrm{N}}$ is an ascent direction, we further employ an Armijo backtracking line search; see [46, Algorithm 3.2]. Specifically, we seek the smallest integer $m\geq 0$ such that $t_{k}=\rho^{m}$ satisfies
@@
where $\rho=0.5$ and $c=10^{-4}$. Note that $\left\langle\Omega_{k}^{\mathrm{N}},\mathrm{g}_{k}\right\rangle>0$, which guarantees that the above procedure terminates after finitely many iterations; see [46, Lemma 3.1]. Consequently, the cost value increases monotonically. Finally, a complete description of the proposed algorithm is given in Algorithm˜2.
This modified Newton method with line search ensures convergence from any initial point (i.e., it enjoys the global convergence property of first-order gradient methods), while achieving quadratic convergence once the iterate is sufficiently close to the solution. Intuitively, during the early iterations, Algorithm˜2 behaves similarly to Algorithm˜1. When the iterate is sufficiently close to the solution (so that $q_{k}>0.5$ and no direction modification is required), the Armijo condition is always satisfied with $t_{k}=1$; see [46, Theorem 3.6 & 3.8]. Consequently, Algorithm˜2 reduces to the standard Newton method.
@@ V Numerical experiments
RGA is shown in [31] to have complexity $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$. Fixing the accuracy $\varepsilon$ and setting $M=1$, this reduces to the well-known quadratic speedup $\mathcal{O}(\sqrt{N})$, i.e., the iteration count grows linearly with $\sqrt{N}$. The proposed RMN is a second-order optimization method, whose improvement targets the dependence on the accuracy $\varepsilon$ rather than on the problem size NN [^3] Optimization literature typically fixes the problem size and studies how to achieve higher accuracy with fewer iterations.. It is therefore expected to retain the same scaling $\mathcal{O}(\sqrt{N})$ with respect to the problem size. However, due to its quadratic convergence, the overall complexity improves to $\mathcal{O}(\sqrt{N}\log\log(1/\varepsilon)).$
@@ VI Discussion
In this work, we develop a deeper understanding of the quantum search problem through the lens of the manifold optimization framework by introducing the Grover-compatible Riemannian modified Newton (RMN) method. While the previous first-order Riemannian gradient ascent (RGA) method [31] successfully recasts quantum search as a maximization problem on the unitary group and achieves $\mathcal{O}(\sqrt{N/M}\log(1/\varepsilon))$ query complexity, it is limited to a linear convergence rate with respect to the target accuracy $\varepsilon$. By incorporating second-order geometric information, we overcome this limitation.
The cornerstone of our RMN method is the theoretical observation that, for projector Hamiltonians $H=H^{2}$, the Riemannian gradient is an eigenvector of the corresponding Riemannian Hessian. This property eliminates the need for costly Hessian inversion, reducing the Newton update to a scaled gradient step. Consequently, our RMN method can achieve a quadratic convergence rate, driving the total complexity down to $\mathcal{O}(\sqrt{N/M}\log\log(1/\varepsilon))$ without incurring additional overhead per iteration. Furthermore, we demonstrate that the classical simulability of the parameter update process is also preserved, allowing for the efficient precomputation of gate angles prior to execution on quantum hardware.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China under the grant numbers 12501419, 12331010 and 12288101, the National Key R&D Program of China under the grant number 2024YFA1012901, the Quantum Science and Technology-National Science and Technology Major Project via Project 2024ZD0301900, and the Fundamental Research Funds for the Central Universities, Peking University. We thank Zongqi Wan for their valuable feedback on the manuscript.
@@ A.4 Riemannian Hessian
A derivation of this Hessian expression can be found in [32, Proposition 1]. To verify the correctness of Eq.˜60, we will apply the second-order Taylor expansion on $\mathrm{U}(N)$. For $\eta=\Omega U\in T_{U}$, using the Riemannian exponential retraction $\mathrm{R}_{U}(\eta)=e^{\Omega}U$ in Eq.˜54, we obtain